# JWST Observations of Starbursts: PAHs Closely Trace the Cool Phase of M82’s Galactic Wind

### v1

@@ Abstract
Stellar feedback drives multiphase gas outflows from starburst galaxies, but the interpretation of dust emission in these winds remains uncertain. To investigate this, we analyze new JWST mid-infrared images tracing polycyclic aromatic hydrocarbon (PAH) emission at 7.7 and 11.3 m from the outflow of the prototypical starburst M82 out to kpc. We find that PAH emission shows significant correlations with CO, H, and X-ray emission within the outflow, though the strengths and behaviors of these correlations vary with gas phase and distance from the starburst. PAH emission correlates strongly with cold molecular gas, with PAH–CO scaling relations in the wind nearly identical to those in galaxy disks despite the very different conditions. The H-PAH correlation indicates that H traces the surfaces of PAH-bearing clouds, consistent with arising from ionized layers produced by shocks. Meanwhile the PAH–X-ray correlation disappears once distance effects are controlled for past 2 kpc, suggesting that PAHs are decoupled from the hot gas and the global correlation merely reflects the large-scale structure of the outflow. The PAH-to-neutral gas ratio remains nearly flat to 2 kpc, with variations following changes in the radiation field. This implies that the product of PAH abundance and dust-to-gas ratio does not change significantly over the inner portion of the outflow. Together, these results demonstrate that PAHs robustly trace the cold phase of M82’s wind, surviving well beyond the starburst and providing a powerful, high-resolution proxy for mapping the life cycle of entrained cold material in galactic outflows.
1 Introduction
Galactic winds are a ubiquitous feature of starburst galaxies. They are driven by stellar feedback and play a vital role in redistributing metals, injecting energy and momentum into the circumgalactic medium (CGM), and regulating star formation (Tumlinson et al., 2017; Veilleux et al., 2020). These outflows are multiphase in nature, spanning large ranges of temperature, densities, and ionization states (Heckman et al., 1990; Veilleux et al., 2005). The hot K X-ray emitting phase, traces the supernova-heated, fast moving, gas that drives the wind (e.g., see Strickland & Heckman, 2009; Lopez et al., 2020; Nguyen & Thompson, 2021). A warm K ionized gas phase is visible in optical line emission (e.g., Xu et al., 2023; Lopez et al., 2025), and can trace the interactions between the hot wind and colder material.
@@ 1 Introduction
Emission from dust grains in the infrared offers a promising direction to explore the physics of cold outflows because the grains are mixed with the outflowing gas (e.g., Engelbracht et al., 2006; Roussel et al., 2010, in M82). With the launch of the James Webb Space Telescope (JWST), emission from polycyclic aromatic hydrocarbons (PAHs) has become a particularly powerful tool to study outflowing material at unprecedented angular resolution and sensitivity (Bolatto et al., 2024; Fisher et al., 2025; Villanueva et al., 2025; Sutter et al., 2025). These small grains produce spectral features at 3.3, 6.2, 7.7, 8.6, 11.2, 12.7, and 17.3 m via their stretching and bending modes (e.g., Puget & Leger, 1989; Tielens, 2008). The illuminating radiation field, the abundance and ionization of PAHs, and dust-to-gas ratio also influence the intensity of PAH emission (e.g., see Leroy et al., 2023; Chown et al., 2025).
In the star-forming disks of spiral galaxies, PAH emission correlates well with CO emission tracing molecular gas at 100 pc resolution (e.g., Leroy et al., 2023; Sandstrom et al., 2023; Chown et al., 2025, and references therein). If a similar correlation holds in galactic outflows, JWST imaging can constrain the morphology, mass, and evolution of cold material as it moves away from the starburst without the need for expensive CO mapping. The fine-scale structure of cold gas is expected to bear the imprint of cloud launching, formation, and destruction, and of mass loading into the hot wind (e.g., see Thompson & Heckman, 2024; Richie & Schneider, 2025).
@@
2 Data
2.1 PAH Emission from JWST/MIRI
We trace 7.7 and 11.3 um PAH emission using JWST images that cover the M82 outflow in the F770W and F1130W filters (doi:https://doi.org/10.17909/ew85-bc51 (catalog 10.17909/ew85-bc51)) from Bolatto et al. (2024) (project GO 1701) which have previously appeared in Fisher et al. (2025) and Villanueva et al. (2025). Following Cronin et al. (in preparation), we subtract a median background estimated from the outer part of the mosaic and confirm the estimate with the JWST Background Calculator111https://github.com/spacetelescope/jwst_backgrounds. We verified that after this, the overall background level of the image appears consistent with previous wide-area Spitzer imaging (Engelbracht et al., 2006; Dale et al., 2009) to MJy sr-1. Also following Cronin et al. (in preparation), a stellar continuum was subtracted from the images using Spitzer IRS data. These images are shown in Figure 1.
2.2 Ancillary High-Resolution Multiwavelength Observations
@@
The CO(1–0) data include both IRAM NOEMA interferometric and 30-m single dish data, and we used the CASA task feather to combine the two datasets. Then we convolved the data to have a round synthesized beam (2.2′′) and estimated the noise ( MJy/sr) based on signal-free regions of the data cube. We produced integrated intensity (moment 0) and associated uncertainty maps by integrating within a three-dimensional mask that includes all position-position-velocity locations where the HI 21-cm map from Martini et al. (2018), the previous IRAM 30-m CO(2–1) map (Leroy et al., 2015), or the Krieger et al. (2021) 30-m CO(1–0) show high significance emission at resolution. This yields a high completeness map of CO(1–0) at resolution (Figure 2). Hereafter we use CO(1–0) and CO interchangeably unless specified (i.e., CO(2–1)).
@@
The Chandra X-ray data were presented by Lopez et al. (2020), using nine archival observations totaling 534 ks of exposures and cover the keV band. The observations were astrometrically aligned, merged, exposure-corrected, and point sources were removed. We matched the JWST, HST H, and Chandra images to the resolution of the CO data following the procedures in Williams et al. (2024).
2.3 Ancillary Low-Resolution Multiwavelength Observations
@@
Because the multiwavelength data have lower angular resolutions than JWST, we make versions of the data at common resolutions of (including JWST, H, CO, and X-ray), (including Spitzer), and (including all data). In each case we reproject the data onto a shared astrometric grid.
3 Results and Discussion
3.1 PAH-Gas Phase Correlation
@@
Our correlation analysis parallels that of Chown et al. (2025), who compared PAH emission to CO(2–1) in galaxy disks. We compare to their results for the disks of galaxies, i.e. omitting galaxy centers. Their best-fit relation, shown in dark blue in the top row of Figure 3, closely resembles what we find for the region of M82 covered by JWST in black. The overlap is excellent near the M82 disk (red points). At larger vertical heights (yellow and green), especially in the F770W filter, the M82 outflow is modestly brighter in CO at fixed PAH intensity than galaxy disks. This may be largely explained by a changing CO(2–1)-to-(1–0) ratio () in the wind222The figure adopts a constant (e.g., Leroy et al., 2022), which is a reasonable first-order approximation for the Chown et al. (2025) sample. Future analysis should account for the variable in both their sample and in the M82 wind..
We also quantify the scatter near and away from the starburst (above/below 750 pc) to test the robustness of the PAH-CO relationship. As shown in Figure 3, the scatter increases away from the starburst. This widening envelope is visually apparent in the green points at large vertical distances in Figure 3, which reflect both the lower signal-to-noise in the CO data (with noise-dominated measurements shown by the blue triangles) and a real increase in intrinsic scatter with height. Assuming a Gaussian distribution, we run a Monte Carlo calculation adopting the power-law fit of CO vs. F770W as the true relation and adding the appropriate statistical noise to each measurement. We find that the scatter is mostly intrinsic333We calculate the intrinsic scatter as . We find and 0.27 dex for the F770W-CO comparison at low and high elevations, respectively. with only moderate statistical scatter increase with height (). The results are nearly identical for F1130W.
@@
While a correlation between CO and PAH emission is expected at these spatial resolutions (Bolatto et al., 2024; Fisher et al., 2025; Villanueva et al., 2025), such close quantitative agreement between our results and Chown et al. (2025) is surprising. The PAH intensity should depend on the strength and hardness of the illuminating radiation field, the dust-to-gas ratio, and the abundance and physical state of the PAHs. These parameters might all be expected to change under the wide range of conditions found in the M82 outflow, as evidenced by the spatial variations in PAH band ratios reported by Beirão et al. (2015), and to differ between the disks of galaxies and the M82 starburst and outflow. We explore the radiation field dependence in Section 3.3.
@@
The PAH-X-ray correlation is puzzling given that small dust grains like PAHs are not expected to survive exposure to the hot wind, as their sputtering times can be one to two orders-of-magnitude shorter than the dynamical time (Richie & Schneider, 2025). The correlation may reflect that a large fraction of the soft X-ray emission is produced by hot gas interacting with cooler material via charge-exchange emission. Such emission has been found to be common in studies of X-ray emitting winds (Okon et al., 2024). In this case, while PAHs cannot survive on their own in the hot gas, if they are mixed with cooler material, their survival times are extended and interactions with the hot gas contribute to the charge-exchange emission.
@@
The H emission is associated with cooler ionized gas and likely represents an intermediate case between CO and X-ray emission. As discussed by Lopez et al. (2025) and Fisher et al. (2025, for Pa), much of the H emission appears to arise from the surfaces of cold gas clouds that are ionized by shocks in the outflow (Shopbell & Bland-Hawthorn, 1998; Beirão et al., 2015), though some may instead be photoionized closer to the starburst. The PAH and H emission also display strikingly similar morphologies (Figure 2), suggesting a common origin. In Appendix B, we present a detailed view of one such structure (cloud 12; Lopez et al. 2025), where the H emission lies along the wind-facing surfaces of the cold cloud traced by PAHs. This configuration highlights where the hot wind may be shocking colder material, and relatedly, producing the observed charge-exchange emission in the soft X-rays. Together, these results support the interpretation that both PAH and H emission originate from the outer layers of cold clouds entrained in the hot outflow or located along the edges of the outflow cone.
@@
We note that the coloring of points in Figure 3 and the images in Figure 2 show that all bands decrease in intensity with increasing distance from the burst along the outflow. Therefore it is possible the correlations reflect a shared dependence of PAH and X-ray intensity on distance from the starburst, not any physical relationship between PAH emission and hot gas. Projection effects may contribute to this effect, as the typical biconical model for M82’s outflow has a volume-filling hot phase with the cooler phases at the cone surface that would appear to overlap (e.g., Leroy et al., 2015; Nguyen & Thompson, 2022).
@@ 3.2 Where are the PAHs?
The middle and right panels also show a complex picture for X-ray and H emission. In both cases, PAHs are bright relative to the other band (red) in the starburst and the galaxy disk (likely due to extinction) and are moderately brighter (salmon) along the edges of the outflow cone, while the other tracers are brighter inside the wind itself. The enhanced PAH emission outside the outflow cone lines up well with the expected multiphase model of galactic winds, where the core is diffuse and hot and the outskirts are cold and dense (Thompson & Heckman, 2024). However, enhanced PAH emission is located within the outflow cone, appearing as patchy filaments. As mentioned in Section 3.1, these locations could be where PAHs are mixed with cooler gas where clouds are entrained in the hot outflow. These patches are less common in the PAH–H ratio map because the two bands exhibit nearly identical morphologies, as shown in Figures 1 and 2 and illustrated in Appendix B for an individual cloud.
3.3 PAH Intensity, Gas Mass, and the Radiation Field

### v2

@@ Abstract
Stellar feedback drives multiphase gas outflows from starburst galaxies, but the interpretation of dust emission in these winds remains uncertain. To investigate this, we analyze new JWST mid-infrared images tracing polycyclic aromatic hydrocarbon (PAH) emission at 7.7 and 11.3 m from the outflow of the prototypical starburst M82 out to kpc. We find that PAH emission shows significant correlations with CO, H, and X-ray emission within the outflow, though the strengths and behaviors of these correlations vary with gas phase and distance from the starburst. PAH emission correlates strongly with cold molecular gas, with PAH–CO scaling relations in the wind nearly identical to those in galaxy disks despite the very different conditions. The H–PAH correlation indicates that H traces the surfaces of PAH-bearing clouds, consistent with arising from ionized layers produced by shocks. Meanwhile the PAH–X-ray correlation disappears once distance effects are controlled for past 2 kpc, suggesting that PAHs are decoupled from the hot gas and the global correlation merely reflects the large-scale structure of the outflow. The PAH-to-neutral gas ratio remains nearly flat to 2 kpc, with variations following changes in the radiation field. This implies that the product of PAH abundance and dust-to-gas ratio does not change significantly over the inner portion of the outflow. Together, these results demonstrate that PAHs robustly trace the cold phase of M82’s wind, surviving well beyond the starburst and providing a powerful, high-resolution proxy for mapping the life cycle of entrained cold material in galactic outflows.
I Introduction
Galactic winds are a ubiquitous feature of starburst galaxies. They are driven by stellar feedback and play a vital role in redistributing metals, injecting energy and momentum into the circumgalactic medium (CGM), and regulating star formation (Tumlinson et al., 2017; Veilleux et al., 2020). These outflows are multiphase in nature, spanning large ranges of temperature, densities, and ionization states (Heckman et al., 1990; Veilleux et al., 2005). The hot K X-ray emitting phase, traces the supernova-heated, fast moving, gas that drives the wind (e.g., see Strickland and Heckman, 2009; Lopez et al., 2020; Nguyen and Thompson, 2021). A warm K ionized gas phase is visible in optical line emission (e.g., Xu et al., 2023; Lopez et al., 2025), and can trace the interactions between the hot wind and colder material.
@@ I Introduction
Emission from dust grains in the infrared offers a promising direction to explore the physics of cold outflows because the grains are mixed with the outflowing gas (e.g., Engelbracht et al., 2006; Roussel et al., 2010, in M82). With the launch of the James Webb Space Telescope (JWST), emission from polycyclic aromatic hydrocarbons (PAHs) has become a particularly powerful tool to study outflowing material at unprecedented angular resolution and sensitivity (Bolatto et al., 2024; Fisher et al., 2025; Villanueva et al., 2025; Sutter et al., 2025). These small grains produce spectral features at 3.3, 6.2, 7.7, 8.6, 11.2, 12.7, and 17.3 m via their stretching and bending modes (e.g., Puget and Leger, 1989; Tielens, 2008). The illuminating radiation field, the abundance and ionization of PAHs, and dust-to-gas ratio also influence the intensity of PAH emission (e.g., see Leroy et al., 2023; Chown et al., 2025).
In the prototypical starburst M82, previous high-resolution mid-infrared spectral mapping has already provided important insight into the behavior of PAHs in outflows. Using IRS integral-field spectroscopy of the central region, Beirão et al. (2008) showed that variations in PAH ratios can be driven by extinction and changes in the local radiation field. Subsequent IRS mapping of the full superwind revealed widespread PAH emission throughout the halo and strong spatial variations in grain size and ionization state (Beirão et al., 2015). These maps demonstrated that H2/PAH ratios in extraplanar gas are enhanced, indicating that shocks contribute significantly to the excitation of the warm molecular component. The PAH 11.3/7.7 ratios were also found to increase with distance from the disk, pointing to larger and more ionized grains in the northern outflow.
In the star-forming disks of spiral galaxies, PAH emission correlates well with CO emission tracing molecular gas at 100 pc resolution (e.g., Leroy et al., 2023; Sandstrom et al., 2023; Chown et al., 2025, and references therein). If a similar correlation holds in galactic outflows, JWST imaging can constrain the morphology, mass, and evolution of cold material as it moves away from the starburst without the need for expensive CO mapping. The fine-scale structure of cold gas is expected to bear the imprint of cloud launching, formation, and destruction, and of mass loading into the hot wind (e.g., see Thompson and Heckman, 2024; Richie and Schneider, 2025).
@@
II Data
II.1 PAH Emission from JWST/MIRI
We trace 7.7 and 11.3 um PAH emission using JWST images that cover the M82 outflow in the F770W and F1130W filters (doi:https://doi.org/10.17909/ew85-bc51 (catalog 10.17909/ew85-bc51)) from Bolatto et al. (2024) (project GO 1701) which have previously appeared in Fisher et al. (2025) and Villanueva et al. (2025). Following Cronin et al. (in preparation), we subtract a median background estimated from the outer part of the mosaic and confirm the estimate with the JWST Background Calculator111https://github.com/spacetelescope/jwst_backgrounds. We verified that after this, the overall background level of the image appears consistent with previous wide-area Spitzer imaging (Engelbracht et al., 2006; Dale et al., 2009) to MJy sr-1. We also account for the stellar and dust continuum that contribute to these PAH bands. Cronin et al. (in preparation) will provide the full details of the continuum subtraction, but in brief, we used the Spitzer IRS spectral cubes of M82 with PAHFIT (Smith et al., 2007) to decompose each spectrum into its emission and continuum components. The resulting continuum maps were then regridded onto the JWST images, and a pixel-by-pixel continuum subtraction was applied. The IRS-derived continuum maps overlap most of the JWST mosaic, but do not cover the full JWST field of view. Where IRS and JWST overlap, we apply the pixel-by-pixel continuum subtraction using the regridded IRS continuum maps. In regions outside the IRS footprint, we adopt a constant continuum fraction equal to the median measured within the IRS coverage for that morphological component. We find that the median continuum contribution to the F770W image is in the northern outflow, in the disk, and in the southern outflow. For the F1130W image, the corresponding values are , , and . We define the disk as regions within pc of the center of M82, with the remaining area corresponding to the northern and southern outflows.
II.2 Ancillary High-Resolution Multiwavelength Observations
@@
The CO(1–0) data include both IRAM NOEMA interferometric and 30-m single dish data, and we used the CASA (CASA Team et al., 2022) task feather to combine the two datasets. Then we convolved the data to have a round synthesized beam (2.2′′) and estimated the noise ( MJy/sr) based on signal-free regions of the data cube. We produced integrated intensity (moment 0) and associated uncertainty maps by integrating within a three-dimensional mask that includes all position-position-velocity locations where the HI 21-cm map from Martini et al. (2018), the previous IRAM 30-m CO(2–1) map (Leroy et al., 2015), or the Krieger et al. (2021) 30-m CO(1–0) show high significance emission at resolution. This yields a high completeness map of CO(1–0) at resolution (Figure 2). Hereafter we use CO(1–0) and CO interchangeably unless specified (i.e., CO(2–1)).
@@
The Chandra X-ray data were presented by Lopez et al. (2020), using nine archival observations totaling 534 ks of exposures and cover the keV band (doi:https://doi.org/10.25574/cdc.320 (catalog 10.25574/cdc.320) ). The observations were astrometrically aligned, merged, exposure-corrected, and point sources were removed. We matched the JWST, HST H, and Chandra images to the resolution of the CO data following the procedures in Williams et al. (2024).
II.3 Ancillary Low-Resolution Multiwavelength Observations
@@
Because the multiwavelength data have lower angular resolutions than JWST, we make versions of the data at common resolutions of (including JWST, H, CO, and X-ray), (including Spitzer), and (including all data). In each case we convolve the data to the coarser PSF and reproject the data onto a shared astrometric grid.
III Results and Discussion
III.1 PAH-Gas Phase Correlation
@@
Our correlation analysis parallels that of Chown et al. (2025), who compared PAH emission to CO(2–1) in galaxy disks. We compare to their results for the disks of galaxies, i.e. omitting galaxy centers. Their best-fit relation, shown in dark blue in the top row of Figure 3, closely resembles what we find for the region of M82 covered by JWST in black. The overlap is excellent near the M82 disk (red points). At larger vertical heights (light and dark red), especially in the F770W filter, the M82 outflow is modestly brighter in CO at fixed PAH intensity than galaxy disks. This may be largely explained by a changing CO(2–1)-to-(1–0) ratio () in the wind222The figure adopts a constant (e.g., Leroy et al., 2022), which is a reasonable first-order approximation for the Chown et al. (2025) sample. Future analysis should account for the variable in both their sample and in the M82 wind..
We also consider a second possibility for the enhanced CO at fixed PAH intensity, which is PAH destruction in the outflow. Beirão et al. (2015) reported elevated H2/PAH ratios in M82’s wind from Spitzer IRS spectra, which could be interpreted as evidence for shocks that both enhance H2 emission and destroy PAHs. The characteristic shock velocities inferred by Beirão et al. (2015) are relatively low ( km s-1). This is below the regime where rapid, efficient PAH destruction is expected. However, Micelotta et al. (2010) show that PAHs can still undergo non-negligible structural damage and erosion for shocks of order km s-1, with increasingly severe destruction at higher velocities. Therefore, while we do not expect shock-driven PAH destruction to be the dominant driver of the CO enhancement relative to the galaxy-disk relation, the Beirão et al. (2015) results imply that modest PAH processing in the outflow may contribute at some level to the observed trends.
We also quantify the scatter near and away from the starburst (above/below 750 pc) to test the robustness of the PAH-CO relationship. We adopt 750 pc as the boundary between the disk and the outflow for three reasons: first, the HI map used in our analysis goes into absorption within the inner 500 pc, leaving the central region undefined (see Fig. 5). Second, when we later convolve the data to a coarser resolution for the vertical profiles, the resulting PSF corresponds to roughly 500 pc, and choosing a larger cutoff avoids contamination from the bright central disk. Finally, visual inspection shows that beyond pc the disk no longer dominates the emission, and the observed signal is primarily from the outflow. As shown in Figure 3, the scatter increases away from the starburst. This widening envelope is visually apparent in the red points at large vertical distances in Figure 3, which reflect both the lower signal-to-noise in the CO data (with noise-dominated measurements shown by the green triangles) and a real increase in intrinsic scatter with height. Assuming a Gaussian distribution, we run a Monte Carlo calculation adopting the power-law fit of CO vs. F770W as the true relation and adding the appropriate statistical noise to each measurement. We find that the scatter is mostly intrinsic333We calculate the intrinsic scatter as . We find and 0.27 dex for the F770W-CO comparison at low and high elevations, respectively. with only moderate statistical scatter increase with height (). The results are nearly identical for F1130W.
@@
While a correlation between CO and PAH emission is expected at these spatial resolutions (Bolatto et al., 2024; Fisher et al., 2025; Villanueva et al., 2025), such close quantitative agreement between our results and Chown et al. (2025) is surprising. The PAH intensity should depend on the strength and hardness of the illuminating radiation field, the dust-to-gas ratio, and the abundance and physical state of the PAHs. These parameters might all be expected to change under the wide range of conditions found in the M82 outflow, as evidenced by the spatial variations in PAH band ratios reported by Beirão et al. (2015), and to differ between the disks of galaxies and the M82 starburst and outflow. We explore the radiation field dependence in Section III.3.
@@
The PAH-X-ray correlation is puzzling given that small dust grains like PAHs are not expected to survive exposure to the hot wind, as their sputtering times can be one to two orders-of-magnitude shorter than the dynamical time (Richie and Schneider, 2025). The correlation may reflect that a large fraction of the soft X-ray emission is produced by hot gas interacting with cooler material via charge-exchange emission. Such emission has been found to be common in studies of X-ray emitting winds (Okon et al., 2024). In this case, while PAHs cannot survive on their own in the hot gas, if they are mixed with cooler material, their survival times are extended and interactions with the hot gas contribute to the charge-exchange emission.
@@
The H emission is associated with cooler ionized gas and likely represents an intermediate case between CO and X-ray emission. As discussed by Lopez et al. (2025) and Fisher et al. (2025, for Pa), much of the H emission appears to arise from the surfaces of cold gas clouds that are ionized by shocks in the outflow (Shopbell and Bland-Hawthorn, 1998; Beirão et al., 2015), though some may instead be photoionized closer to the starburst. The PAH and H emission also display strikingly similar morphologies (Figure 2), suggesting a common origin. In Appendix B, we present a detailed view of one such structure (cloud 12; Lopez et al. 2025), where the H emission lies along the wind-facing surfaces of the cold cloud traced by PAHs. This configuration highlights where the hot wind may be shocking colder material, and relatedly, producing the observed charge-exchange emission in the soft X-rays. Together, these results support the interpretation that both PAH and H emission originate from the outer layers of cold clouds entrained in the hot outflow or located along the edges of the outflow cone.
@@
We note that the coloring of points in Figure 3 and the images in Figure 2 show that all bands decrease in intensity with increasing distance from the burst along the outflow. Therefore it is possible the correlations reflect a shared dependence of PAH and X-ray intensity on distance from the starburst, not any physical relationship between PAH emission and hot gas. Projection effects may contribute to this effect, as the typical biconical model for M82’s outflow has a volume-filling hot phase with the cooler phases at the cone surface that would appear to overlap (e.g., Leroy et al., 2015; Nguyen and Thompson, 2022).
@@ III.2 Where are the PAHs?
The middle and right panels also show a complex picture for X-ray and H emission. In both cases, PAHs are bright relative to the other band (red) in the starburst and the galaxy disk and are moderately brighter (salmon) along the edges of the outflow cone, while the other tracers are brighter inside the wind itself. The enhanced PAH emission compared to the ionized gas phases in the disk likely arises from a combination of effects. As shown by the H image in Figure 2, prominent dust lanes are present in the disk of M82. This extinction suppresses the observed H emission. A similar effect can attenuate the softer X-ray photons ( keV) that contribute significantly to the broad-band Chandra image, leading to an apparent enhancement of PAH emission relative to the ionized gas in the disk.
The PAH/[Ne II] ratio is similar to PAH/H (Ne I ionization potential is 21.6 ev) but less affected by extinction due to Ne II’s wavelength. This ratio has been measured by Beirão et al. (2008) and Beirão et al. (2015) for M82’s disk and outflow, respectively. A comparison of these studies shows that PAH/[Ne II] decreases by approximately an order of magnitude from the disk (PAH/[Ne II]) to the outflow (PAH/[Ne II]). This indicates a real physical change beyond only extinction. The higher ratio in the dist could reflect large amounts of diffuse PAH emission along the line of sight through the edge-on disk. Alternatively, PAH heating conditions may differ between the disk and the outflow, with a more intense and harder radiation field in the disk. Finally, the absorption of ionizing photons in dusty H II regions in the dense disk may suppress the [Ne II] emission.
It is also possible that PAH processing occurs within the outflow. Although the shock velocities inferred in M82’s wind by Beirão et al. (2015) are relatively low, such shocks can still produce non-negligible processing of PAHs (Micelotta et al., 2010), which may modestly reduce their emission. In addition, the radiation field decreases with increasing distance from the disk, and in Section III.3 we show that this decline closely tracks the decrease in the gas-weighted PAH emission. These changes in the physical environment and in PAH properties can therefore contribute to the observed variations in the enhanced PAH-to-gas ratio maps. A full analysis of the composition of PAHs in M82’s wind will be presented in Cronin et al. (in preparation).
@@
The enhanced PAH emission observed outside the outflow cone is consistent with the expected multiphase structure of galactic winds, in which the core is hot and diffuse while the outskirts are cooler and denser (Thompson and Heckman, 2024). Interestingly, enhanced PAH emission is also located within the outflow cone, appearing as patchy filaments. As mentioned in Section III.1, these locations could be where PAHs are mixed with cooler gas where clouds are entrained in the hot outflow. These patches are less common in the PAH–H ratio map because the two bands exhibit nearly identical morphologies, as shown in Figures 1 and 2 and illustrated in Appendix B for an individual cloud.
III.3 PAH Intensity, Gas Mass, and the Radiation Field